OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (anonymous):

@ricky15 you know thats not you in your icon right?

OpenStudy (anonymous):

that is me

OpenStudy (anonymous):

@ricky15 its a tumblr boy

OpenStudy (anonymous):

-x - 2y = -4 -x=2y-4 x=-2y+4 so D

OpenStudy (anonymous):

@Juliette32801 your comments do not have anything to do with the question.

OpenStudy (anonymous):

@warningiflirt so....

OpenStudy (anonymous):

@ricky15 if you want I can send another picture of "you"

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

@ricky15 ok

OpenStudy (anonymous):

@ricky15 when i find it I will

OpenStudy (anonymous):

sure